INAF queue observing, 2018 June 30-July 01

Observer: F. Cusano, R. Speziali
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson (on call)
Instrument: LUCI

Summary

At the sunset the sky is clear

Night duration (12\xB0 el twilight) 4:10- 10:35 UT = 6 h 25m

Observing time: 5h 05min (74%)

Weather time loss: h m (%)

technical time loss 1h 20m*(26%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibration phz1119 LUCI 1+2 dark NDIT=1, DIT=120.0 MER, NORM   dark for Polletta phz1119
Calibration telluric LUCI 1+2 dark NDIT=6, DIT=5.0 LIR INT   dark for Polletta telluric
Calibration G016.58 LUCI 1+2 dark NDIT=30, DIT=10.0 LIR INT   dark for Moscadelli H2
Calibration G016.58 LUCI 1+2 dark NDIT=10, DIT=3.0 LIR INT   dark for Moscadelli Ks
Moscadelli G016.58 LUCI 1 H2 1.1h   completed
Standard FS 140 LUCI 1 H2, Ks      
Moscadelli G012.91 LUCI 1 H2, Ks 1.25h, 10min?    
Caccianiga gb6j2357 MODS 1 red grating 1.5h   completed
Standard Feige 110 MODS 1 red grating      

Details

00:32 starting dark LUCI1+2 for Polletta phz1119
  • luci1.20180701.0001-0005
  • luci2.20180701.0001-0005

01:00 dark for telluric
  • luci1.20180701.0006-0010
  • luci2.20180701.0006-0010

1:10 dark for Moscadelli G016.58, for H2 filter
  • luci1.20180701.0011-0015
  • luci2.20180701.0011-0015

1:46 dark for Moscadelli G016.58, for Ks filter
  • luci1.20180701.0016-0020
  • luci2.20180701.0016-0020

3:33 sending Rossi 130803A, spec
  • 3:34 preset failed, the to and Doug are checking, problem in the pointing
  • 3:35 error in the mask, no mask used
  • 3:40 problem with pointing solved, but the mask is in unknown position
  • 3:50 D. Thompson is working on it

4:09 starting again Rossi
  • 4:17 dimm seeing 0.8"
  • 4:24 mask error again, can not move mask in FPU
  • can not solve the problem moving to Moscadelli program LUCi1 only

4:45 slewing to standard star of Moscadelli, FS 140, LUCI1 only
  • 4:52 problem in collimation, the TO is using the idl script and sending correction manually, not good we add a guide star
  • 4:57 sending again the script with a guide star
  • 5:08 starting taking data filter H2
  • 5:14 starting Ks

5:19 slewing to G016.58, H2 filter
  • 5:21 no guide star found, the to check the pointing
  • 5:23 sending again the script
  • 5:26 starting H2 filter
  • seeing 0.8" on first image
  • dimm seeing 1.2" at 6:08
  • dimm seeing increasing to 1.5" at 6:12, measured on imaging 1.1"
  • dimm seeing 1.1" at 6:21

6:46 slewing to G12.91, H2 filter
  • 7:07 dimm seeing 0.8"
  • 7:32 seeing 1.1" on guider

07:54 starting G12.91 Ks filter
  • lost 15 minutes for collimation
  • 60 not good
  • 8:24 repeating Ks (we prefer to stay on Ks, otherwise the preset of H2 will cancel the collimation)

8:34 starting again H2 G12.91
  • 8:35 dimm seein 0.8"
  • 8:37 starting integration
  • dimm seeing 0.7" at 8:48

8:55 switching to MODS1

9:18 slewing to Caccianiga, gb6j2357, acqMODS gb6j2357_mods1red.acq.txt
  • acq mods1.20180701.0002-004
  • 9:25 spec mods1.20180701.0005-0010
  • dimm seeing 1.1" at 9:39
  • 10:16 sending again the spec script
  • seeing 0.8" on guider at 10:12

11:08* slewing to standard Feige 110
  • spec mods1.20180701.0013-0015
Topic revision: r6 - 06 Jul 2018, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback